OpenStudy (anonymous):

What is the value of x in the equation 4x + 8y = 40, when y = 0.8? 4.6 8.4 10 12

OpenStudy (anonymous):

4x + 8y = 40 4x + 8(0.8) = 40 What's 8 * 0.8?

OpenStudy (anonymous):

6.4

OpenStudy (anonymous):

Yes, so we have: 4x + 6.4 = 60 Now subtract 6.4 to both sides, what's 60 - 6.4?

OpenStudy (anonymous):

53.6

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now divide 53.6 / 4

OpenStudy (anonymous):

Because we have: 4x = 53.6

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i got 13.4

OpenStudy (anonymous):

Yes, 13.4 is correct.